OpenStudy (anonymous):

Find all solutions in the interval [0, 2pi). (sin x) (cos x) = 0

OpenStudy (raden):

set sinx = 0 or cosx=0 this is a start, do u know what angles to get sinx=0, cosx=0 ?

OpenStudy (anonymous):

sin =0 and cos=pi/2

OpenStudy (raden):

sinx=0, it satisfied for x=0, x=pi, x=2pi now, find the solution for cosx=0 ?

OpenStudy (anonymous):

pi/2, 3pi/2?

OpenStudy (raden):

opss.. nevermind for x=2pi, it didnt including in interval [0,2pi) :) yes, u are right

OpenStudy (anonymous):

Thank you! do you think you can help me with my last one? its kind of tricky?

OpenStudy (raden):

post it

OpenStudy (anonymous):

on here? or close this one and make another?

OpenStudy (raden):

just here

OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (raden):

for this time u must fator out, first... 7 sin2x - 14 sin x + 2 = -5 or 7 sin2x - 14 sin x + 2 + 5 = 0 7 sin2x - 14 sin x + 7= 0 simplify, divided by 7 gives sin^2 x - 2sinx + 1 = 0 now, can u factor out this

OpenStudy (anonymous):

and what happens to the 7?

OpenStudy (raden):

if the equation of 7 sin^2 x - 14 sin x + 7= 0 divided by 7 to both sides, it means 7/7 sin^2 x - 14/7 sin x + 7/7= 0/7 u get sin^2 x - 2sin x + 1= 0 right ?

OpenStudy (anonymous):

ok you can't just factor it out?

OpenStudy (raden):

i can, can u ? :)

OpenStudy (anonymous):

yea....? :/ but if you factor it out, what happens to the 7?

OpenStudy (anonymous):

so x= pi/2, 3pi/2?

OpenStudy (raden):

u dont need the 7 again, it passed away :)

OpenStudy (anonymous):

ok is my top answer correct? :)

OpenStudy (raden):

finally, the answer will be x=pi/2 only do u know why ?

OpenStudy (anonymous):

no....?

OpenStudy (anonymous):

why not 3pi/2?

OpenStudy (raden):

look this: from sin^2 x - 2sin x + 1= 0, factor out gives (sinx - 1)(sinx - 1)=0 it satisfied for sinx - 1 = 0, solve for x

OpenStudy (anonymous):

sinx=1 so it equals pi/2 oh because it is not asking for a negative? and 3pi/2=-1?

OpenStudy (raden):

yeah, u are right :)

OpenStudy (anonymous):

oh ok! thank you!!

OpenStudy (raden):

very welcome :)